# Spectral and Temporal Denoising for Differentially Private Optimization

### v1

@@ 3.3 Frequency-Domain Denoising
$\mathcal{G}_{\rho}(z):=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\odot\mathcal{F}(z)\bigr{)},\qquad z\in\mathbb{R}^{d}.$
This filtering step yields the estimate $\widehat{g}_{t}=\mathcal{G}_{\rho}(g_{t})$. Since $\mathcal{G}_{\rho}$ is a linear operator with spectral mask $\Phi_{\rho}$, this operation has complexity $O(d\log d)$ and does not distort the signal beyond a known attenuation factor. That is, the covariance of $\widehat{g}_{t}$ is a spectrally reweighted version of $\mathrm{Cov}(g_{t})$, which we exploit in the Kalman update below.
@@ 3.4 FFT-Enhanced Kalman Filter
We adopt the scalar-gain Kalman filtering approximation introduced in [32], which simplifies the covariance matrices to scalar multiples of the identity. Specifically, we let $P_{t}=p_{t}I_{d}$, $K_{t}=\kappa I_{d}$, and estimate the Hessian action using a finite-difference formula with hyperparameter $\gamma>0$.
Prediction Step. Given $\tilde{g}_{t-1}$, we predict the next gradient by using a first-order approximation based on finite differences:
$\tilde{g}_{t|t-1}=\tilde{g}_{t-1}+\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\frac{\nabla f(x_{t}+\gamma d_{t-1};\xi)-\nabla f(x_{t};\xi)}{\gamma},$ (P)
where $d_{t-1}:=x_{t}-x_{t-1}=-\eta\tilde{g}_{t-1}$. This approximates the action of the local Hessian without explicitly computing second-order derivatives.
@@
where $\kappa\in(0,1)$ is the Kalman gain that balances the reliance on the prediction versus the new (denoised) observation. This form ensures that the update direction incorporates temporal consistency across iterations while attenuating the influence of high-frequency noise.
Together, Eqs. (P) and (C) constitute a computationally lightweight Kalman filtering mechanism enhanced by frequency-domain denoising. It achieves $O(d\log d)$ per-step complexity and enhances DP optimization without sacrificing privacy guarantees.
@@
1:initial point $x_{0}$, base optimiser $\mathsf{Opt}$, learning rate $\eta$, gain $\kappa$, FD parameter $\gamma$, high–frequency ratio $\rho$, clipping bound $C$, noise scale $\sigma_{w}$.
2:$\tilde{g}_{-1}\leftarrow 0,\;d_{-1}\leftarrow 0$
@@
5:     Compute shaped DP gradient
$g_{t}\leftarrow\frac{1}{B}\!\!\sum_{\xi\in\mathcal{B}_{t}}\!\!\operatorname{clip}\Bigl{(}\nabla f(x_{t};\xi),\,C\Bigr{)}+\tilde{w}_{t},\quad\tilde{w}_{t}\sim\mathcal{N}\bigl{(}0,\sigma_{w}^{2}I_{d}\bigr{)}$
6:     $\widehat{g}_{t}\leftarrow\mathcal{G}_{\rho}(g_{t})$triangleright FFT denoising
7:     $\tilde{g}_{t|t-1}\leftarrow$ Eq. (P)
@@ 3.5 Additional Discussion
The high-frequency shaping in Eq. (3) intentionally pushes privacy noise into spectral regions that matter least for optimization. Because the Kalman filter relies on low-frequency temporal correlations captured by Eqs. (P)–(C), the FFT step removes most of the injected disturbance before the gain $\kappa$ is applied, resulting in a provably lower steady-state covariance.
@@
$\Sigma_{\tilde{w}}=\mathcal{F}^{-1}\cdot\Phi_{\rho}^{2}\cdot\mathcal{F}\cdot\Sigma_{w}\cdot\mathcal{F}^{-1}\cdot\Phi_{\rho}^{2}\cdot\mathcal{F},$
@@
3.5.1 Connections to DiSK
Scalar–gain Kalman simplification.
Our FFT-Enhanced Kalman Filter (FFTKF) inherits the scalar–gain reduction of DiSK [32], wherein both the state covariance $P_{t}$ and the Kalman gain $K_{t}$ are isotropic:
$P_{t}=p_{t}I_{d},\quad K_{t}=\kappa I_{d}.$
@@ 3.5.1 Connections to DiSK
$\boldsymbol{H}_{t}d_{t-1}\approx\frac{\nabla F(x_{t}+\gamma d_{t-1})-\nabla F(x_{t})}{\gamma},$
@@
FFT-based noise shaping.
While DiSK performs time-domain exponential smoothing, FFTKF additionally reshapes the injected DP noise to concentrate its energy in the high-frequency spectrum:
$\tilde{w}_{t}:=\mathcal{F}^{-1}\bigl{(}\Phi_{\rho}\odot\mathcal{F}(w_{t})\bigr{)},\quad(\Phi_{\rho})_{k}=\begin{cases}1,&k<k_{0},\\[2.0pt] 1-\rho e^{-\alpha(k-k_{0})},&k\geq k_{0},\end{cases}$
@@
Computational footprint.
Compared with DPSGD, FFTKF requires:
  1. 1.
    one additional forward pass per iteration to compute the finite-difference directional gradient, and
  2. 2.
    two in-place 1-D FFTs: a forward transform $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$.
Both operations scale as $O(d\log d)$, while the state vector $\tilde{g}_{t}$ and difference direction $d_{t}$ are stored as $O(d)$ vectors. Thus, FFTKF matches the memory profile of DiSK [32] but enables more precise noise attenuation with marginal overhead.
Privacy guarantee.
Since the FFT operation is orthonormal, it preserves the $\ell_{2}$ norm:
$\|\tilde{w}_{t}\|_{2}=\|\Phi_{\rho}\odot\mathcal{F}(w_{t})\|_{2}\leq\|w_{t}\|_{2}.$
Thus, FFT-based reshaping does not increase the sensitivity of the privatized quantity. The overall privacy budget $(\varepsilon,\delta)$ remains exactly that of DPSGD and DiSK, guaranteed by
@@ 4 Theoretical Analysis: Privacy-Utility Trade-off
Let the FFT operator be $\mathcal{F}:\mathbb{R}^{d}\!\to\!\mathbb{C}^{d}$ with inverse $\mathcal{F}^{-1}$, as introduced in Section 3.1. Fix a pivot index $k_{0}=\lfloor\lambda d\rfloor$ ($\lambda\!\in\!(0,1)$) and a high–frequency attenuation ratio $\rho\in(0,1)$. Define the diagonal spectral mask
$\Phi_{\rho}=\operatorname{diag}\!\bigl{(}\underbrace{1,\dots,1}_{k_{0}},\underbrace{1-\rho,\dots,1-\rho}_{d-k_{0}}\bigr{)},$
and the deterministic post-processing map $P(g)=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\,\mathcal{F}g\bigr{)}.$ Given a privatised gradient $g_{t}$, the filtered release is $\hat{g}_{t}:=P(g_{t}).$
Privacy is preserved.
Proposition 4.1 (Post-processing invariance).
Because $P$ is data-independent, $\hat{g}_{t}$ is $(\varepsilon,\delta)$-DP whenever the DiSK gradient $g_{t}$ is $(\varepsilon,\delta)$-DP.This follows from the post-processing theorem of differential privacy [8, Thm. 2.1].
Sensitivity note. The mask satisfies $\|\Phi_{\rho}\|_{2}=1$; hence $P$ does not increase the $\ell_{2}$-sensitivity of its input. The Gaussian noise scale $\sigma_{w}$ chosen for DiSK therefore continues to satisfy the target $(\varepsilon,\delta)$ budget. Consequently, Algorithm 1 inherits exactly the same overall $(\varepsilon,\delta)$ guarantee as standard DP-SGD/DiSK, computed with the moments accountant over $T$ iterations.
Bias and covariance.
Lemma 4.2 (Effect of the low-pass mask).
Write $g_{t}=\nabla F(x_{t})+\eta_{t}$ with $\eta_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$. Let $\rho^{\star}=\bigl{(}k_{0}+(1-\rho)^{2}(d-k_{0})\bigr{)}/d$. Then
@@
Proof.
Unitary invariance of $\mathcal{F}$ yields the stated mean and covariance. Eigenvalues of $A$ are $1$ (multiplicity $k_{0}$) and $1-\rho$ (multiplicity $d-k_{0}$). ∎
Remark.  “Bias” in Lemma 4.2 refers to $E[\hat{g}_{t}]-\nabla F(x_{t})$; filtering does not introduce systematic noise bias but scales the signal by $A$.
Updated convergence bound.
Lemma 4.2 replaces the isotropic noise term $d\sigma_{w}^{2}$ in the DiSK analysis with $\rho^{\star}d\sigma_{w}^{2}$ and introduces a multiplicative bias factor $1-\rho$. Repeating the steps of Zhang el yields:
Theorem 4.3 (Privacy–utility with FFT filtering).
Under Assumptions A1–A3 and the same $(\eta,\kappa,\gamma)$ schedule as in Zhang et al., Algorithm 1 satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(x_{t})\|^{2}$ $\leq\frac{2\bigl{(}F(x_{0})-F^{\star}+\beta\|\nabla F(x_{0})\|^{2}\bigr{)}}{C_{1}\eta T}$
$+\frac{2(\beta+\eta^{2}L)\kappa^{2}}{C_{1}\eta}\!\Bigl{[}(2+|1+\gamma|)\rho^{\star}\,d\sigma_{w}^{2}+\tfrac{\sigma_{\!SGD}^{2}}{B}\Bigr{]}$
$+\rho^{2}\,G_{T},$
@@
$C_{1}=(1+\kappa-2\eta L)-4(\beta+\eta^{2}L)(1-\kappa)^{2}L^{2}\eta(2+|1+\gamma|).$
Practical choice and independence of the mask.
In all experiments we fix $\lambda=\tfrac{1}{2}$ and $\rho=0.5$ a priori (i.e. independently of any individual training sample); this gives $\rho^{\star}=0.625$ and $\rho^{2}=0.25$. Thus the DP-noise contribution is reduced by $37.5\%$ while the extra bias inflates the optimization term by at most $25\%$, yielding a provably tighter trade-off than plain DiSK.
@@ 5 Experimental Results
In this section, we explore how the FFT-Enhanced Kalman Filter (FFTKF) improves the performance of differential privacy (DP) optimizers on various models, datasets, and privacy budgets. The utilization of FFT for the purpose of reshaping the DP noise in the frequency domain is undertaken with the objective of preserving the essential low-frequency gradient signal, while concomitantly directing privacy noise into spectral regions. For code and implementation details, please refer to \$ https://github.com/OpenNN-Theory/KalmanDP-BEDB.
@@ 5.1 Experimental Settings
Model Method CIFAR10 [13] CIFAR100 [13] MNIST [16] Tiny-ImgNet [15] CNN5 [14] DPAdam [27] 60.83 23.34 91.18 10.93 DISK [32] 70.57 39.62 92.52 23.45 FFTKF (ours) 70.86 40.80 92.62 22.62 WRN-16 [31] DPAdam [27] 53.81 15.53 91.18 10.63 DISK[32] 69.33 35.19 92.52 24.13 FFTKF (ours) 70.38 35.91 92.62 24.61 WRN-28 [31] DPAdam [27] 53.72 14.74 91.18 9.94 DISK[32] 71.22 37.79 92.52 26.96 FFTKF (ours) 72.58 38.93 92.62 27.63 WRN-40 [31] DPAdam [27] 54.50 14.05 91.18 8.92 DISK [32] 72.13 37.31 92.52 27.41 FFTKF (ours) 73.73 37.95 92.62 27.83 ViT-small [7] DPAdam [27] 50.98 19.83 91.18 13.47 DISK [32] 58.79 32.44 92.52 25.70 FFTKF (ours) 59.85 32.46 92.62 25.96
Table 1: Test accuracy (%) under $(\epsilon=4)$ across four datasets and five model architectures. The proposed FFTKF method consistently improves or matches the performance of DPAdam and DISK baselines, particularly on CIFAR and Tiny-ImageNet benchmarks.

### v2

@@ 3.3 Frequency-Domain Denoising
$\mathcal{G}_{\rho}(z)$ $:=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\odot\mathcal{F}(z)\bigr{)},\qquad z\in\mathbb{R}^{d}.$ (19)
This filtering step yields the estimate $\widehat{g}_{t}=\mathcal{G}_{\rho}(g_{t})$. Since $\mathcal{G}_{\rho}$ is a linear operator with spectral mask $\Phi_{\rho}$, this operation has complexity $O(d\log d)$ and does not distort the signal beyond a known attenuation factor. The covariance of $\widehat{g}_{t}$ is a spectrally reweighted version of $\mathrm{Cov}(g_{t})$, which we exploit in the Kalman update below [12].
@@ 3.4 FFT-Enhanced Kalman Filter
We adopt the scalar-gain Kalman filtering approximation introduced in [32], which simplifies the covariance matrices to scalar multiples of the identity. Specifically, we let $P_{t}=p_{t}I_{d}$, $K_{t}=\kappa I_{d}$, and estimate the Hessian action using a privatized finite-difference formula with hyperparameter $\gamma>0$.
Prediction Step. Given $\tilde{g}_{t-1}$, we predict the next gradient by using a first-order approximation based on privatized finite differences:
$\tilde{g}_{t|t-1}=$ $\tilde{g}_{t-1}$ (20)
$+\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\frac{\text{clip}(\nabla f(x_{t}+\gamma d_{t-1};\xi),C)}{\gamma}$ (21)
$-\sum_{\xi\in\mathcal{B}_{t}}\frac{\text{clip}(\nabla f(x_{t};\xi),C)}{\gamma}+w_{t}^{\text{fd}},$ (22)
where $d_{t-1}:=x_{t}-x_{t-1}=-\eta\tilde{g}_{t-1}$, $w_{t}^{\text{fd}}\sim\mathcal{N}(0,\sigma_{\text{fd}}^{2}I_{d})$ is additional noise for privacy, and clipping is applied to bound sensitivity. This approximates the action of the local Hessian without explicitly computing second-order derivatives.
@@
where $\kappa\in(0,1)$ is the Kalman gain that balances the reliance on the prediction versus the new (denoised) observation. This form ensures that the update direction incorporates temporal consistency across iterations while attenuating the influence of high-frequency noise. Together, Eqs. (22) and (C) constitute a computationally lightweight Kalman filtering mechanism enhanced by frequency-domain denoising. The per-step complexity is $O(d\log d)$ for FFT operations plus $O(d)$ for two gradient evaluations and finite-difference computation, achieving overall efficiency while enhancing DP optimization without sacrificing privacy guarantees.
@@
1:initial point $x_{0}$, base optimiser $\mathsf{Opt}$, learning rate $\eta$, gain $\kappa$, FD parameter $\gamma$, high–frequency ratio $\rho$, clipping bound $C$, noise scales $\sigma_{w}$, $\sigma_{\text{fd}}$.
[…]
@@
5:  Compute privatized gradient with isotropic noise
$g_{t}\leftarrow\frac{1}{B}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}\bigl{(}\nabla f(x_{t};\xi),C\bigr{)}+w_{t},\quad w_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$ (23)
6:  $\widehat{g}_{t}\leftarrow\mathcal{G}_{\rho}(g_{t})$ $\triangleright$ FFT denoising
7:  $\tilde{g}_{t|t-1}\leftarrow$ Eq. (22) $\triangleright$ Privileged finite-difference prediction
@@ 3.5 Additional Discussion
The high-frequency shaping in Eq. (17) intentionally pushes privacy noise into spectral regions that matter least for optimization. Because the Kalman filter relies on low-frequency temporal correlations captured by Eqs. (22)–(C), the FFT step removes most of the injected disturbance before the gain $\kappa$ is applied, resulting in a provably lower steady-state covariance.
@@
$\Sigma_{\tilde{w}}=\mathcal{F}^{-1}\cdot\Phi_{\rho}^{2}\cdot\mathcal{F}\cdot\Sigma_{w}\cdot\mathcal{F}^{-1}\cdot\Phi_{\rho}^{2}\cdot\mathcal{F},$ (24)
@@
Scalar–gain Kalman simplification. Our FFT-Enhanced Kalman Filter (FFTKF) inherits the scalar–gain reduction of DiSK [32], wherein both the state covariance $P_{t}$ and the Kalman gain $K_{t}$ are isotropic:
$P_{t}=p_{t}I_{d},\quad K_{t}=\kappa I_{d}.$ (25)
@@
$\boldsymbol{H}_{t}d_{t-1}\approx\frac{\nabla F(x_{t}+\gamma d_{t-1})-\nabla F(x_{t})}{\gamma},$ (26)
@@
FFT-based noise shaping. While DiSK performs time-domain exponential smoothing, FFTKF additionally reshapes the injected DP noise to concentrate its energy in the high-frequency spectrum:
$\begin{aligned} \tilde{w}_{t}&=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\odot\mathcal{F}(w_{t})\bigr{)},\end{aligned}\text{ }\begin{aligned} (\Phi_{\rho})_{k}&=\begin{cases}1,&k<k_{0},\\[4.0pt] 1-\rho e^{-\alpha(k-k_{0})},&k\geq k_{0}.\end{cases}\end{aligned}$ (27)
@@
Computational footprint. Compared with DPSGD, FFTKF requires one additional forward pass per iteration to compute the finite-difference directional gradient, a forward transform $\mathcal{F}$, and its inverse $\mathcal{F}^{-1}$. Both operations scale as $O(d\log d)$, while the state vector $\tilde{g}_{t}$ and difference direction $d_{t}$ are stored as $O(d)$ vectors. Thus, FFTKF matches the memory profile of DiSK [32] but enables more precise noise attenuation with marginal overhead.
Privacy guarantee. Since the FFT operation is orthonormal, it preserves the $\ell_{2}$ norm:
$\|\tilde{w}_{t}\|_{2}=\|\Phi_{\rho}\odot\mathcal{F}(w_{t})\|_{2}\leq\|w_{t}\|_{2}.$ (28)
Thus, FFT-based reshaping does not increase the sensitivity of the privatized quantity. The overall privacy budget $(\varepsilon,\delta)$ remains exactly that of DPSGD and DiSK, guaranteed by the post-processing property of differential privacy, which ensures that any transformation applied after the privatization step cannot degrade the original privacy guarantees.
@@ 4 Theoretical Analysis: Privacy-Utility Trade-off
We theoretically analyze our method based on KF-filter method for differential privacy [20].
Let the FFT operator be $\mathcal{F}:\mathbb{R}^{d}\!\to\!\mathbb{C}^{d}$ with inverse $\mathcal{F}^{-1}$, as introduced in Section 3.1. Fix a pivot index $k_{0}=\lfloor\lambda d\rfloor$ ($\lambda\!\in\!(0,1)$) and a high–frequency attenuation ratio $\rho\in(0,1)$. Define the diagonal spectral mask:
$\Phi_{\rho}=\operatorname{diag}\!\bigl{(}\underbrace{1,\dots,1}_{k_{0}},\underbrace{1-\rho,\dots,1-\rho}_{d-k_{0}}\bigr{)},$ (29)
and the deterministic post-processing map $P(g)=\mathcal{F}^{-1}\!\bigl{(}\Phi_{\rho}\,\mathcal{F}g\bigr{)}.$ Given a privatised gradient $g_{t}$, the filtered release is $\hat{g}_{t}:=P(g_{t}).$ Then, privacy is preserved.
Proposition 1: Post-processing invariance Because $P$ is data independent, $\hat{g}_{t}$ is $(\varepsilon,\delta)$-DP whenever the DiSK gradient $g_{t}$ is $(\varepsilon,\delta)$-DP.This follows from the post-processing theorem of differential privacy [8, Thm. 2.1].
The mask satisfies $\|\Phi_{\rho}\|_{2}=1$; hence $P$ does not increase the $\ell_{2}$-sensitivity of its input. The Gaussian noise scale $\sigma_{w}$ chosen for DiSK therefore continues to satisfy the target $(\varepsilon,\delta)$ budget. Consequently, Algorithm 1 inherits exactly the same overall $(\varepsilon,\delta)$ guarantee as standard DP-SGD/DiSK, computed with the moments accountant over $T$ iterations.
Lemma 1: Effect of the low-pass mask Write $g_{t}=\nabla F(x_{t})+\eta_{t}$ with $\eta_{t}\sim\mathcal{N}(0,\sigma_{w}^{2}I_{d})$. Let $\rho^{\star}=\bigl{(}k_{0}+(1-\rho)^{2}(d-k_{0})\bigr{)}/d$. Then
@@
Proof. Unitary invariance of $\mathcal{F}$ yields the stated mean and covariance. Eigenvalues of $A$ are $1$ (multiplicity $k_{0}$) and $1-\rho$ (multiplicity $d-k_{0}$).
Remark.  “Bias” in Lemma 4 refers to $E[\hat{g}_{t}]-\nabla F(x_{t})$; filtering does not introduce systematic noise bias but scales the signal by $A$.
Lemma 4 replaces the isotropic noise term $d\sigma_{w}^{2}$ in the DiSK analysis with $\rho^{\star}d\sigma_{w}^{2}$ and introduces a multiplicative bias factor $1-\rho$. Repeating the steps of Theorem 2 [32] yields:
Theorem 2. Privacy–utility with FFT filtering Under Assumptions A1–A3 and the same $(\eta,\kappa,\gamma)$ schedule as in Algorithm 1 satisfies
$\frac{1}{T}$ $\sum_{t=0}^{T-1}\mathbb{E}\|\nabla F(x_{t})\|^{2}$ (30)
$\leq\frac{2\bigl{(}F(x_{0})-F^{\star}+\beta\|\nabla F(x_{0})\|^{2}\bigr{)}}{C_{1}\eta T}$ (31)
$\qquad+\frac{2(\beta+\eta^{2}L)\kappa^{2}}{C_{1}\eta}\!\Bigl{[}(2+|1+\gamma|)\rho^{\star}\,d\sigma_{w}^{2}+\tfrac{\sigma_{\!SGD}^{2}}{B}\Bigr{]}$ (32)
$\qquad+\rho^{2}\,G_{T},$ (33)
@@
$C_{1}=(1+\kappa-2\eta L)-4(\beta+\eta^{2}L)(1-\kappa)^{2}L^{2}\eta(2+|1+\gamma|).$ (34)
Practical choice and independence of the mask. In all experiments we fix $\lambda=\tfrac{1}{2}$ and $\rho=0.5$ a priori (i.e. independently of any individual training sample); this gives $\rho^{\star}=0.625$ and $\rho^{2}=0.25$. Thus the DP-noise contribution is reduced by $37.5\%$ while the extra bias inflates the optimization term by at most $25\%$, yielding a provably tighter trade-off than plain DiSK.
@@ 5 Experimental Results
Model Method CIFAR10 CIFAR100 MNIST Tiny-ImgNet
CNN5 DPAdam 60.83 23.34 91.18 10.93
DISK 70.57 39.62 92.52 23.45
FFTKF 70.86 40.80 92.62 22.62
WRN-16 DPAdam 53.81 15.53 91.18 10.63
DISK 69.33 35.19 92.52 24.13
FFTKF 70.38 35.91 92.62 24.61
WRN-28 DPAdam 53.72 14.74 91.18 9.94
DISK 71.22 37.79 92.52 26.96
FFTKF 72.58 38.93 92.62 27.63
WRN-40 DPAdam 54.50 14.05 91.18 8.92
DISK 72.13 37.31 92.52 27.41
FFTKF 73.73 37.95 92.62 27.83
ViT-small DPAdam 50.98 19.83 91.18 13.47
DISK 58.79 32.44 92.52 25.70
FFTKF 59.85 32.46 92.62 25.96
Figure 3: Left: Test accuracy (%) under $(\epsilon=4)$ across four datasets and five model architectures. Right: Test accuracy across $(\rho,\epsilon)$ at epoch 40.
In this section, we explore how the FFT-Enhanced Kalman Filter (FFTKF) improves the performance of differential privacy (DP) optimizers on various models, datasets, and privacy budgets. The utilization of FFT for the purpose of reshaping the DP noise in the frequency domain is undertaken with the objective of preserving the essential low-frequency gradient signal, while concomitantly directing privacy noise into spectral regions.